OpenStudy (anonymous):

In radian measure, the angles of a 45-45-90 triangle are

OpenStudy (anonymous):

pi=180

OpenStudy (anonymous):

it is asking for radian measure a 45 - 45 - 90 triangle convert to radians and ge t \[\frac{\pi}{4},\frac{\pi}{4},\frac{\pi}{2}\]

OpenStudy (anonymous):

@satellite73 can you help me with my question?

OpenStudy (anonymous):

to convert, multiply by \(\pi\) and divide by \(180\) so for example \[60\]degrees becomes \[\frac{60\pi}{180}=\frac{\pi}{3}\]

OpenStudy (anonymous):

@GABI_J if i can i will

OpenStudy (anonymous):

After this question?